Find the surface whose tangent plane cut off an intercept of constant length k from the axis z


1
Expert's answer
2021-07-26T06:46:49-0400

Let "f(x,y,z)=0" be the equation of the surface, "P(x_P, y_P, z_P)" - arbitrary point on it.

The equation of the tangent plane at "P" is

"\frac{\partial f}{\partial x}(P)(x-x_P) + \frac{\partial f}{\partial y}(P)(y-y_P)+\frac{\partial f}{\partial z}(P)(z-z_P)=0"


This plane intersects the axis z at the point "Q=(0,0,z_Q)" , and by the condition "z_Q=k" is independent on "P" .

Since "Q" lies on the tangent plane, then

"\frac{\partial f}{\partial x}(P)(-x_P) + \frac{\partial f}{\partial y}(P)(-y_P)+\frac{\partial f}{\partial z}(P)(k-z_P)=0"

Consider cylindrical coordinates "(r, \theta, z)" . Then "x=r\cos\theta, y=r\sin\theta" ,

"\frac{\partial}{\partial r}(x,y,z)=\frac{\partial}{\partial r}(r\cos\theta, r\sin\theta, z)=(\cos\theta, \sin\theta, 0) = (\frac{x}{r}, \frac{y}{r}, 0)" .

Hence

"\frac{\partial}{\partial r}=\frac{x}{r}\frac{\partial}{\partial x} + \frac{y}{r}\frac{\partial}{\partial y}"

"x \frac{\partial}{\partial x} + y \frac{\partial}{\partial y}=r \frac{\partial}{\partial r}"


Therefore, we can rewrite the obtained equation as follows:

"\frac{\partial f}{\partial z}(P)(k-z_P)=\frac{\partial f}{\partial x}(P)x_P + \frac{\partial f}{\partial y}(P)y_P =r \frac{\partial f}{\partial r}(P)"


The first partial solution to this equation is, evidently, the coordinate function "\theta".

However, it is a multi-valued function on "\mathbb{R}^3 \setminus Oz" . So, we should take "f_1 = e^{i\theta}" instead of it.


The second partial solution can be found of the form "f_2 = z + g(r)", where "g(r)" is unknown function.

"r \frac{\partial f_2}{\partial r} = rg'(r)"

"k \frac{\partial f_2}{\partial r} = k"

Hence "rg'(r)=k" and "g(r)= k\log r" (partial solution).


The functions "f_1" and "f_2" are functionally independent, therefore the general solution can be written as

"f(x,y,z) = \Phi(f_1, f_2)= \Phi(e^{i\theta}, z+k\log r)" , where "\Phi" is arbitrary smooth function on "\mathbb{R}^3 \setminus Oz" .


Answer. "f(r, \theta, z) = \Phi(f_1, f_2)= \Phi(e^{i\theta}, z+k\log r)" (in cylindrical coordinates)